OpenStudy (anonymous):

Which of the following expressions could give the total amount of a $45 restaurant bill with a 15% tip added on?

OpenStudy (anonymous):

51.75?

terenzreignz (terenzreignz):

Okay, this is tricky... if you have choices, get 15% of that choice and ADD it to that choice... if you get 45, then that's the one :P

terenzreignz (terenzreignz):

And it can't possibly be more than 45 :D

OpenStudy (anonymous):

my choices (0.15)($45) (1.15)($45) 1 + (0.15)($45) $45 + 0.15

OpenStudy (anonymous):

last one i susppose

terenzreignz (terenzreignz):

Oh, I must have misunderstood your question... I'm sorry you were right the first time :)

OpenStudy (anonymous):

ill still give you a medal for efort

OpenStudy (texaschic101):

it would be 45 + (.15(45)....would that be the 2nd or the 3rd one

OpenStudy (anonymous):

2nd

terenzreignz (terenzreignz):

45 + (0.15)(45) = 45(1+0.15) = 45(1.15) etc :3

OpenStudy (texaschic101):

got it....thanks

OpenStudy (anonymous):

so it the 3rd choice

OpenStudy (texaschic101):

thats what terenzreignz said

terenzreignz (terenzreignz):

No, it's the second choice...

OpenStudy (texaschic101):

no he said 2nd...sorry

terenzreignz (terenzreignz):

in the end, I got (45)(1.15) which is the same as (1.15)(45)

OpenStudy (anonymous):

lol thanks